# REOBench: Benchmarking Robustness of Earth Observation Foundation Models

### v1

@@ 3.6 Visual Question Anaswering
Table 5 reports VQA performance across various image perturbations. Following the VRSBench protocol [26], VQA performance is evaluated using the GPT-4-based matching accuracy [^2] We use the gpt-4o-mini-2024-07-18 model to compute the matching accuracy for VQA.. From Table 5, it is evident that all LLM-based models experience a moderate decline in performance under image perturbations. Overall, VHM [52] achieves the best accuracy across both clean and noisy images. LHRS-Bot [50], RS-LLaVA [51], and Falcon [54], despite showing relatively lower overall accuracy, exhibit less sensitivity to image corruptions. Additionally, the GeoChat [12] fine-tuned on the VRSBench training set surpasses zero-shot models in terms of absolute performance, while also demonstrating robustness to image perturbations.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\Delta_{\text{TP}}$
@@
GeoChat [12] ViT-L 56.63 53.89 54.82 55.14 55.99 55.88 55.44 56.22 54.08 54.04 55.06 1.57
LHRS-Bot [50] ViT-L 35.72 35.72 35.69 35.72 35.72 35.72 35.72 35.72 35.34 35.72 35.56 0.16
SkySenseGPT [53] ViT-L 60.21 59.26 59.73 57.93 59.64 59.21 58.27 59.63 59.17 57.27 58.90 1.31
VHM [52] ViT-L 61.72 60.91 61.07 60.40 61.49 60.91 60.91 61.12 59.97 60.39 60.90 0.82
RS-LLaVA [51] ViT-L 57.25 57.04 57.14 55.45 57.25 57.14 55.97 57.21 55.25 55.82 56.47 0.78
Falcon [54] DaViT-B 33.27 32.83 32.70 32.19 33.30 33.43 32.85 32.76 32.97 31.55 32.73 0.53
GeoChatft [12] ViT-L 75.79 75.13 74.97 73.84 75.63 74.89 74.46 75.43 74.76 72.77 74.65 1.14
@@ 3.7 Visual Grounding
Table 6 presents the zero-shot visual grounding performance of comparing methods. We report grounding accuracy at an IoU threshold of 0.5. As shown in the upper part of Table 6, all methods experience noticeable declines in performance under image perturbations. The GeoGround [59] model achieves the best performance on both clean and perturbed images, with a grounding accuracy of 75.93%. Under image perturbations, its performance declines modestly, with an average drop of 3.4%. In comparison, the Falcon [54] model, despite not being trained on VRSBench, demonstrates competitive visual grounding capability, but experiences a more significant degradation in performance when exposed to image corruptions. The fine-tuned GeoChat [12] model shows substantial improvements over its zero-shot counterpart; however, it also experiences a noticeable decline in performance under noisy conditions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\Delta_{\text{TP}}$
@@
GeoChat [12] ViT-L 18.96 17.09 16.54 16.52 16.19 16.61 16.93 17.09 16.91 16.57 16.72 2.24
VHM [52] ViT-L 37.20 34.66 35.29 34.18 35.48 35.01 35.78 35.54 32.21 34.58 34.74 2.46
GeoGround [59] ViT-L 75.93 73.57 71.57 71.30 72.23 73.23 72.92 74.06 72.11 71.77 72.53 3.40
Falcon [54] DaViT-B 73.30 71.31 69.92 65.83 68.61 70.79 64.28 71.04 68.17 59.53 67.72 5.58
GeoChatft [12] ViT-L 55.50 53.79 52.20 50.51 53.06 53.11 51.57 54.23 52.99 49.82 52.36 3.14
@@ 4 Discussion
As shown in Fig. 4, vision-centric foundation models (MIM-based) tend to suffer greater performance degradation under visual perturbations compared to vision-language models (CL- and CLM-based). This difference is especially pronounced in image-level scene classification tasks, where MIM-based models exhibit an average performance drop exceeding 20%. In contrast, vision-language models consistently demonstrate stronger robustness across tasks, maintaining performance drops below 10% in most cases. This is probably due to the complementary grounding effect of language supervision. We also note that the robustness gap between vision-centric and vision-language models is less significant for segmentation and detection tasks.
@@ 4.2 Robustness Across Different Tasks
Fig. 4 further highlights that vulnerability to perturbations varies substantially across tasks. MIM-based models are particularly sensitive in classification tasks, while CL-based models maintain greater stability across classification, segmentation, and detection tasks. This can be attributed to the contrastive objective, which encourages learning of invariant and robust representations. LLM-based models, on the other hand, show the smallest performance degradation in vision-language tasks such as image captioning, visual question answering (VQA), and phrase grounding—typically below 3%. These results suggest that LLM-based methods excel in corruption-robust generalization, particularly in tasks that benefit from multimodal alignment.
Figure 2: Robustness across different tasks and model architectures. We report the average $\Delta_{TP}$ across models.
Figure 3: Robustness across different backbone sizes. We report the average $\Delta_{TP}$ for RemoteCLIP and GeoRSCLIP.
Figure 4: Robustness across different types of corruptions. We report the average $\Delta_{TP}$ across models.
@@ 4.3 Robustness Across Different Backbone Sizes
We compare the average performance of RemoteCLIP [9] and GeoRSCLIP [10] using ViT-B and ViT-L backbones across scene classification, semantic segmentation, and object detection tasks. As illustrated in Fig. 4, the impact of backbone size varies by task. For scene classification, the larger backbone (ViT-L) demonstrates greater robustness, showing less performance degradation under image corruptions. However, for fine-grained tasks such as semantic segmentation and object detection, ViT-L suffers larger performance drops compared to ViT-B. This suggests that while a larger backbone may enhance robustness in high-level recognition tasks, it may also amplify sensitivity to image corruptions in pixel- or region-level tasks.
@@ 4.4 Robustness Across Perturbation Types
As shown in Fig. 4, the performance degradation of MIM-, CL-, and LLM-based models varies notably across different visual perturbations. Motion blur causes the most severe drop, especially for MIM, which loses over 20% in performance, indicating a high sensitivity to spatial distortions. In contrast, translation has the least impact, suggesting minimal disruption to pattern recognition. Across perturbation types, LLM-based models consistently exhibit the strongest robustness, maintaining performance drops below 5% in most cases. This reinforces the value of language supervision in promoting the learning of more semantic and perturbation-invariant features.
@@ 7 Broader Impact
Supplementary of REOBench: Benchmarking Robustness of Earth Observation Foundation Models
[…]
Chunbo Luo2, Lu Liu2, Mykola Pechenizkiy6, Xiao Xiang Zhu5, Tianjin Huang2,6
1 University of Reading, UK 2 University of Exeter, UK 3 South China Normal University, China
4 Mohamed bin Zayed University of Artificial Intelligence, United Arab Emirates
5 Technical University of Munich, Germany 6 Eindhoven University of Technology, NL
@@ A.2 Data Organization
[…]
.1 root/. .2 AID. .3 AID_train.zip. .3 AID_test.zip. .3 AID_JSON. .2 Potsdam. .3 Potsdam_Images_trian.zip. .3 Potsdam_Anns_trian.zip. .3 Potsdam_Images_test.zip. .3 Potsdam_Anns_test.zip. .2 DIOR. .3 DIOR_Images_trian.zip. .3 DIOR_Anns_trian.zip. .3 DIOR_Images_test. .4 clean.zip; brightness_contrast.zip; cloud.zip; compression_artifacts.zip…. .3 DIOR_Anns_test. .4 clean.zip; rotate.zip; scale.zip; translate.zip. .2 VRSBench. .3 VRSBench_Images_trian.zip. .3 VRSBench_train.json. .3 VRSBench_Images_test. .4 clean.zip; brightness_contrast.zip; cloud.zip; compression_artifacts.zip…. .3 VRSBench_EVAL_Cap.json. .3 VRSBench_EVAL_referring.json. .3 VRSBench_EVAL_vqa.json.
@@ E.2 Image Pertubations
In table 8, we report parameter configures for different noise levels and types. S1 to S5 represent the severity levels of each corruption, with S1 being the mildest and S5 the most severe. The “Parameter” column lists the key control variables for each type of image corruption. For Gaussian noise, $\sigma$ represents the standard deviation that controls the amplitude of the Gaussian noise; for salt-and-pepper noise, the amount specifies the proportion of pixels randomly replaced with black or white; Gaussian blur and motion blur use kernel size to define the size of the convolution kernel; brightness and contrast corruption is controlled using $b/c$ to represent brightness offset and contrast scaling respectively; for the cloud corruption, density denotes the cloud coverage, which is controlled based on the Perlin noise threshold [77]; the parameter for haze simulation is intensity, a blending ratio that determines the degree of mixing with a white layer; data gaps are defined by number/width of stripes, which specifies the number and width of missing regions; Compression Artifacts are governed by the JPEG quality parameter, with lower values indicating stronger compression artifacts. In terms of geometric transformations, rotation is defined by angle, scaling is represented by scale ratio, and translation is described by the maximum offset in displacement. These parameters allow for systematic control and severity grading of each corruption type.
@@ F.1 More Implementation Details
For the semantic segmentation task, all experiments are conducted on a single NVIDIA A100 GPU with 40 GB of memory. The experiments are implemented using the MMSegmentation [80] framework, with UPerHead used as the primary decoder head and FCNHead as the auxiliary head. All training and evaluation are performed on the Potsdam dataset, where each original image ($000\times 6,$ 000 × 6 , 000 pixels) is divided into 144 non-overlapping patches of $512\times 512$ pixels for model training and inference. The resulting dataset comprises 5,472 image patches, split into 3,456 samples for training and 2,016 for testing. Training is performed for 12 epochs using the AdamW optimizer [81], with an initial learning rate of 6e-5, a batch size of 8, and a weight decay of 0.05.
For the object detection task, all experiments are conducted on a single NVIDIA A100 GPU with 80 GB of memory. The experiments are implemented using the MMRotate framework [82], with Oriented R-CNN employed as the detection head. All images in the DIOR-R dataset are resized to $800\times 800$ pixels. The official trainval split is used for training, while the test split is used for both clean evaluation and as the basis for our noise-augmented test set. For fine-tuning, we extracted the last-layer feature map from each model’s pretrained backbone and constructed a four-level feature pyramid via up/down-sampling, which is then fed into a newly initialized Oriented R-CNN head. The detector is trained for 12 epochs using the AdamW optimizer [81], with an initial learning rate of 1e-5, a batch size of 1, and a learning rate decay of 0.1 applied at epochs 8 and 11.

### v2

@@ 3.6 Visual Question Anaswering
Table 5 reports VQA performance across various image perturbations. Following the VRSBench protocol [26], VQA performance is evaluated using the GPT-4-based matching accuracy [^2] We use the gpt-4o-mini-2024-07-18 model to compute the matching accuracy for VQA.. From Table 5, it is evident that all LLM-based models experience a moderate decline in performance under image perturbations. Overall, VHM [52] achieves the best accuracy across both clean and noisy images. LHRS-Bot [50], RS-LLaVA [51], and Falcon [54], despite showing relatively lower overall accuracy, exhibit less sensitivity to image corruptions. Additionally, the GeoChat [12] fine-tuned on the VRSBench training set surpasses zero-shot models in terms of absolute performance, while exhibiting a slightly smaller performance drop under perturbations, indicating improved robustness.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\mathcal{R}_{\text{TP}}$
@@
GeoChat [12] ViT-L 56.63 53.89 54.82 55.14 55.99 55.88 55.44 56.22 54.08 54.04 55.06 2.77
LHRS-Bot [50] ViT-L 35.72 35.72 35.69 35.72 35.72 35.72 35.72 35.72 35.34 35.72 35.56 0.45
SkySenseGPT [53] ViT-L 60.21 59.26 59.73 57.93 59.64 59.21 58.27 59.63 59.17 57.27 58.90 2.18
VHM [52] ViT-L 61.72 60.91 61.07 60.40 61.49 60.91 60.91 61.12 59.97 60.39 60.90 1.33
RS-LLaVA [51] ViT-L 57.25 57.04 57.14 55.45 57.25 57.14 55.97 57.21 55.25 55.82 56.47 1.36
Falcon [54] DaViT-B 33.27 32.83 32.70 32.19 33.30 33.43 32.85 32.76 32.97 31.55 32.73 1.59
GeoChatft [12] ViT-L 75.79 75.13 74.97 73.84 75.63 74.89 74.46 75.43 74.76 72.77 74.65 1.50
@@ 3.7 Visual Grounding
Table 6 presents the zero-shot visual grounding performance of comparing methods. We report grounding accuracy at an IoU threshold of 0.5. As shown in the upper part of Table 6, all methods experience noticeable declines in performance under image perturbations. The GeoGround [59] model achieves the best performance on both clean and perturbed images, with a grounding accuracy of 75.93% and the smallest drop of 4.48%. In comparison, the Falcon [54] model, despite not being trained on VRSBench, demonstrates competitive visual grounding capability, but experiences a more significant degradation in performance when exposed to image corruptions. The fine-tuned GeoChat [12] model shows substantial improvements over its zero-shot counterpart, with a substantially reduced performance drop under noisy conditions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\mathcal{R}_{\text{TP}}$
@@
GeoChat [12] ViT-L 18.96 17.09 16.54 16.52 16.19 16.61 16.93 17.09 16.91 16.57 16.72 11.81
VHM [52] ViT-L 37.20 34.66 35.29 34.18 35.48 35.01 35.78 35.54 32.21 34.58 34.74 6.61
GeoGround [59] ViT-L 75.93 73.57 71.57 71.30 72.23 73.23 72.92 74.06 72.11 71.77 72.53 4.48
Falcon [54] DaViT-B 73.30 71.31 69.92 65.83 68.61 70.79 64.28 71.04 68.17 59.53 67.72 7.61
GeoChatft [12] ViT-L 55.50 53.79 52.20 50.51 53.06 53.11 51.57 54.23 52.99 49.82 52.36 5.66
@@ 4 Discussion
As shown in Fig. 4, vision-centric foundation models (MIM-based) tend to suffer greater performance degradation under visual perturbations compared to vision-language models (CL- and VLM-based). This difference is especially pronounced in image-level scene classification tasks, where MIM-based models exhibit an average performance drop exceeding 25%. In contrast, vision-language models consistently demonstrate stronger robustness across tasks, maintaining performance drops below 10% in most cases. This is probably due to the complementary grounding effect of language supervision. We also note that the robustness gap between vision-centric and vision-language models is less significant for segmentation and detection tasks.
@@ 4.2 Robustness Across Different Tasks
Fig. 4 further highlights that vulnerability to perturbations varies substantially across tasks. MIM-based models are particularly sensitive in classification tasks, while CL-based models maintain greater stability across classification, segmentation, and detection tasks. This can be attributed to the contrastive objective, which encourages learning of invariant and robust representations. LLM-based models, on the other hand, show the smallest performance degradation in vision-language tasks such as image captioning and visual question answering (VQA)—typically below 5%. These results suggest that LLM-based methods excel in corruption-robust generalization, particularly in tasks that benefit from multimodal alignment.
Figure 2: Robustness across different tasks and model architectures. We report the average $\mathcal{R}_{\text{TP}}$ across models.
Figure 3: Robustness across different backbone sizes. We report the average $\mathcal{R}_{\text{TP}}$ for RemoteCLIP and GeoRSCLIP.
Figure 4: Robustness across different types of corruptions. We report the $\mathcal{R}_{\text{TP}}$ across models.
@@ 4.3 Robustness Across Different Backbone Sizes
For scene classification, the larger backbone (ViT-L) demonstrates greater robustness, showing less performance degradation under image corruptions. However, for fine-grained tasks such as semantic segmentation and object detection, ViT-L suffers larger performance drops compared to ViT-B. This suggests that while a larger backbone may enhance robustness in high-level recognition tasks, it may also amplify sensitivity to image corruptions in pixel- or region-level tasks.
@@ 4.4 Robustness Across Perturbation Types
As shown in Fig. 4, the performance degradation of MIM-, CL-, and LLM-based models varies notably across different visual perturbations. Motion blur causes the most severe drop, especially for MIM, which loses around 60% in performance, indicating a high sensitivity to spatial distortions. In contrast, translation has the least impact, suggesting minimal disruption to pattern recognition. Across perturbation types, LLM-based models consistently exhibit the strongest robustness, maintaining performance drops below 10% in most cases. This reinforces the value of language supervision in promoting the learning of more semantic and perturbation-invariant features.
4.5 Robustness Across Compound Perturbations
Real-world images often suffer from multiple simultaneous degradations (e.g., haze combined with noise). To investigate this, we evaluate object detection performance under selected compound perturbations on the DIOR-R dataset. The results from Table 7 reveal that models exhibit substantially lower accuracy under compound perturbations compared to single perturbations. For example, the detection performance of RemoteCLIP drops from 56.72% under Brightness to 40.58% under all three combined perturbations. Moreover, certain combinations of perturbations, such as Brightness + Compression, produce synergistic effects, causing performance declines that exceed the sum of individual perturbation effects.
Table 7: Object Detection performance (mAP) on DIOR-R under compound corruptions
Model Clean Brightness Clouds Compression Brightness Brightness Clouds All Three
+ Clouds + Compression + Compression
Brightness q q q q
Clouds q q q q
Compression q q q q
RemoteCLIP 60.40 56.72 56.28 56.78 49.98 45.36 52.57 40.58
GeoRSCLIP 60.20 56.28 56.04 56.08 50.27 44.90 52.39 40.94
4.6 Robustness in Multispectral Remote Sensing
In addition to high-resolution RGB imagery, multispectral data provide complementary spectral information that can improve scene understanding and robustness in Earth Observation. To assess the robustness of EO foundation models beyond optical imagery, we conduct preliminary experiments on two representative multispectral datasets: fMoW-Sentinel2 [60] and BigEarthNet [61], using two recently proposed foundation models, SatMAE [5] and EarthDial [62]. The results, summarized in Table 8, show substantial performance drops under image perturbations, underscoring the brittleness of current EO foundation models when applied to multispectral data.
Table 8: Scene classification performance on the multispectral dataset across different image perturbations.
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\mathcal{R}_{\text{TP}}$
Contrast Artifacts Gaps Blur Noise Blur Pepper
SatMAE [5] fMoW-S2 [60] 59.75 37.46 58.69 33.58 50.01 29.35 36.64 41.57 40.12 38.93 51.79 43.35 59.68 43.43 27.31
EarthDial [62] BigEarthNet [61] 46.521 31.47 46.48 45.40 34.97 36.37 38.69 33.25 39.84 29.51 22.71 39.18 45.30 36.93 20.62
@@ 7 Broader Impact
Acknowledgments and Disclosure of Funding
This work is supported by the EDF-KOIOS project and used the Dutch national e-infrastructure with the support of the SURF Cooperative using the funding of the projects EINF-12538, EINF-10925 and NWO-2023.027. We would like to express our deepest gratitude to the anonymous reviewers whose insightful comments and suggestions significantly improved the quality of this paper.
@@ A.2 Data Organization
root/
    AID
    AID_train.zip
    AID_test.zip
    AID_JSON
    Potsdam
    Potsdam_Images_trian.zip
    Potsdam_Anns_trian.zip
    Potsdam_Images_test.zip
    Potsdam_Anns_test.zip
    DIOR
    DIOR_Images_trian.zip
    DIOR_Anns_trian.zip
    DIOR_Images_test
    clean.zip; brightness_contrast.zip; cloud.zip; compression_artifacts.zip...
    DIOR_Anns_test
    clean.zip; rotate.zip; scale.zip; translate.zip
    VRSBench
    VRSBench_Images_trian.zip
    VRSBench_train.json
    VRSBench_Images_test
    clean.zip; brightness_contrast.zip; cloud.zip; compression_artifacts.zip...
    VRSBench_EVAL_Cap.json
    VRSBench_EVAL_referring.json
    VRSBench_EVAL_vqa.json
@@ E.2 Image Pertubations
In table 10, we report parameter configures for different noise levels and types. S1 to S5 represent the severity levels of each corruption, with S1 being the mildest and S5 the most severe. The “Parameter” column lists the key control variables for each type of image corruption. For Gaussian noise, $\sigma$ represents the standard deviation that controls the amplitude of the Gaussian noise; for salt-and-pepper noise, the amount specifies the proportion of pixels randomly replaced with black or white; Gaussian blur and motion blur use kernel size to define the size of the convolution kernel; brightness and contrast corruption is controlled using $b/c$ to represent brightness offset and contrast scaling respectively; for the cloud corruption, density denotes the cloud coverage, which is controlled based on the Perlin noise threshold [80]; the parameter for haze simulation is intensity, a blending ratio that determines the degree of mixing with a white layer, and in subsequent visual fidelity experiments we extend the severity levels up to 9 with corresponding intensity values of 0.70, 0.80, 0.90, and 0.95 for S6–S9, respectively; data gaps are defined by number/width of stripes, which specifies the number and width of missing regions; Compression Artifacts are governed by the JPEG quality parameter, with lower values indicating stronger compression artifacts. In terms of geometric transformations, rotation is defined by angle, scaling is represented by scale ratio, and translation is described by the maximum offset in displacement. These parameters allow for systematic control and severity grading of each corruption type.
E.3 Perturbation Fidelity Analysis
To assess the visual fidelity of our synthetic perturbations, we compute the Fréchet Inception Distance (FID) [81] between corrupted images and their corresponding clean counterparts. Specifically, we report FID scores for four benchmarks with different image resolutions: DIOR ($\times800$), VRSBench ($\times512$), Potsdam ($\times512$), and AID ($\times256$). For each dataset, FID was calculated across five severity levels (1–5) of perturbations, using the clean dataset as the reference distribution. Table 11 shows that FID values remain consistently low across severity levels, suggesting that the corrupted datasets are visually similar to real-world imagery and therefore suitable for robustness evaluation.
Table 11: FID scores between corrupted datasets and their corresponding clean datasets across different severity levels. Lower values indicate higher visual similarity.
Severity DIOR ($\times800$) VRSBench ($\times512$) Potsdam ($\times512$) AID ($\times256$)
1 2.36 3.84 6.06 14.07
2 5.46 7.67 10.35 23.60
3 9.67 11.69 17.06 32.37
4 14.21 14.76 24.86 39.31
5 19.81 20.06 37.89 49.03
To evaluate the visual fidelity of our synthetic weather perturbations relative to real-world corruptions, we conduct experiments on two benchmark datasets: SEN12MS-CR [82] for cloud perturbations and RRSHID [83] for haze perturbations. For each dataset, we apply our synthetic perturbations to the clean subsets and compute FID against the corresponding real corrupted subsets, which serve as references for natural weather effects. As shown in Table 12, the FID scores remain below 130 across all severity levels, indicating that the generated perturbations are visually similar to real-world conditions. For cloud perturbations on SEN12MS-CR, the FID increases consistently with severity, suggesting that lower levels more closely resemble natural cloud coverage, whereas higher levels introduce progressively stronger distortions. For haze perturbations on RRSHID, the severity range is extended to level 9. The FID decreases up to level 5 and rises thereafter, implying that moderate haze (around level 5) best matches the distribution of real-world haze samples.
Table 12: Comparative FID scores for synthetic perturbations against real-world weather datasets across different severity levels. Lower values indicate higher visual similarity.
Dataset 1 2 3 4 5 6 7 8 9
SEN12MS-CR (clouds) [82] 105.56 117.27 123.72 125.70 127.03
RRSHID (haze) [83] 107.46 100.79 90.79 80.52 73.23 73.83 88.37 128.83 190.30
@@ F.1 More Implementation Details
For the semantic segmentation task, all experiments are conducted on a single NVIDIA A100 GPU with 40 GB of memory. The experiments are implemented using the MMSegmentation [86] framework, with UPerHead used as the primary decoder head and FCNHead as the auxiliary head. All training and evaluation are performed on the Potsdam dataset, where each original image ($6,000\times 6,000$ pixels) is divided into 144 non-overlapping patches of $512\times 512$ pixels for model training and inference. The resulting dataset comprises 5,472 image patches, split into 3,456 samples for training and 2,016 for testing. Training is performed for 12 epochs using the AdamW optimizer [87], with an initial learning rate of 6e-5, a batch size of 8, and a weight decay of 0.05.
For the object detection task, all experiments are conducted on a single NVIDIA A100 GPU with 80 GB of memory. The experiments are implemented using the MMRotate framework [88], with Oriented R‑CNN employed as the detection head. All images in the DIOR‑R dataset are resized to $800\times 800$ pixels. The official trainval split is used for training, while the test split is used for both clean evaluation and as the basis for our noise-augmented test set. For fine-tuning, we extracted the last-layer feature map from each model’s pretrained backbone and constructed a four-level feature pyramid via up/down-sampling, which is then fed into a newly initialized Oriented R‑CNN head. The detector is trained for 12 epochs using the AdamW optimizer [87], with an initial learning rate of 1e-5, a batch size of 1, and a learning rate decay of 0.1 applied at epochs 8 and 11.